A student prepares 100.00 mL of 0.133 M NaCl in water. 12.5 mL of the solution is then measured with a 25 mL graduated cylinder and transferred to a 50.00 mL volumetric flask which is then filled to the mark with deionized water.

what is the molarity of NaCl in the 50.00 mL volumetric flask?
Answer #1

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 0.133 M

V1 = 12.5 mL

V2 = 50 mL

use:

M1*V1 = M2*V2

M2 = (M1*V1)/V2

M2 = (0.133*12.5)/50

M2 = 0.0333 M

Answer: 0.0333 M